# Point-GN: A Non-Parametric Network Using Gaussian Positional Encoding for Point Cloud Classification

### v1

@@ 4.4 Shape Classification on ScanObjectNN
7070707080808080909090901001001001001101101101101201201201201301301301308080808082828282848484848686868688888888KAccuracy (%)ModelNet40 AverageModelNet40 BestScanObjectNN AverageScanObjectNN Best
18181818272727273636363645454545545454546363636372727272818181819090909099999999Dimension
222233334444Stage
0.20.20.20.20.250.250.250.250.30.30.30.30.350.350.350.350.40.40.40.40.450.450.450.45Sigma
@@ 4.7 Ablation Study
We conducted an ablation study to evaluate the impact of different model configurations on the performance of Point-GN. Specifically, we examined the effect of four key factors: the dimension of Gaussian Positional Encoding (GPE), the number of neighbors $k$ in the k -Nearest Neighbors algorithm, the standard deviation $\sigma$ of the Gaussian function, and the number of stages in the model.
Effect of GPE Dimension. The dimensionality of the Gaussian Positional Encoding (GPE) significantly affects model performance. As shown in Fig. 5, increasing the dimension up to 27 improves accuracy, reaching a peak of 85.3%. However, beyond 45, further increases in dimension result in diminishing returns, with accuracy slightly declining. This suggests that while higher dimensions can capture more complex features, excessively high dimensions add unnecessary complexity without substantial performance gains.
Impact of Sigma ($\sigma$). The $\sigma$ parameter of the Gaussian kernel ($\sigma$) determines the degree of locality in feature aggregation. For the ModelNet40 dataset, $\sigma=0.3$ and $\sigma=0.35$ yield the highest accuracy, effectively preserving both local and broader spatial contexts. On the other hand, For the ScanObjectNN dataset, we find that $\sigma=0.4$ is most effective, achieving the best accuracy by maintaining a balance between capturing fine-grained details and minimizing noise. Smaller values ($\sigma<0.3$) fail to aggregate sufficient local features, while larger values ($\sigma>0.4$) introduce excessive smoothing, negatively impacting performance. These trends are illustrated in Fig. 5.
Effect of Number of Neighbors ($k$). The number of neighbors $k$ used in the k-NN algorithm also plays a critical role in model performance. Our experiments show that $k=120$ offers the best trade-off between computational efficiency and accuracy. Smaller values, such as $k=70$, fail to capture sufficient local context, while larger values, like $k=130$, introduce irrelevant neighbors that confuse the model, hindering its ability to discern fine-grained geometric features. The performance comparison is shown in Fig. 5.
Effect of Number of Stages. Increasing the number of stages in the model generally improves accuracy. The 4-stage configuration achieves the highest accuracy of 85.3%, suggesting that deeper models are better equipped to capture complex spatial relationships and improve classification performance. Refer to Fig. 5 for a detailed comparison of performance across different stage configurations.

### v2

@@ 4.4 Shape Classification on ScanObjectNN
7070707080808080909090901001001001001101101101101201201201201301301301308080808082828282848484848686868688888888(a) KAccuracy (%)ModelNet40 AverageModelNet40 BestScanObjectNN AverageScanObjectNN Best
18181818272727273636363645454545545454546363636372727272818181819090909099999999(b) Dimension
222233334444(c) Stage
0.20.20.20.20.250.250.250.250.30.30.30.30.350.350.350.350.40.40.40.40.450.450.450.45(d) Sigma
@@ 4.7 Ablation Study
We conducted an ablation study to evaluate the impact of different model configurations on the performance of Point-GN. Specifically, we examined the effect of four key factors: the number of neighbors $k$ in the k-Nearest Neighbors algorithm, the dimension of Gaussian Positional Encoding (GPE), the number of stages in the model and the standard deviation $\sigma$ of the Gaussian function.
Effect of Number of Neighbors ($k$). The number of neighbors $k$ used in the k-NN algorithm also plays a critical role in model performance. Our experiments show that $k=\textbf{120}$ offers the best trade-off between computational efficiency and accuracy. Smaller values, such as $k=70$, fail to capture sufficient local context, while larger values, like $k=130$, introduce irrelevant neighbors that confuse the model, hindering its ability to discern fine-grained geometric features. The performance comparison is shown in Fig. 5 (a).
Effect of GPE Dimension. The dimensionality of the Gaussian Positional Encoding (GPE) significantly affects model performance. As shown in Fig. 5 (b), increasing the dimension up to 27 improves accuracy, achieving the best performance in both ModelNet40 and ScanObjectNN datasets. However, beyond 45, further increases in dimension result in diminishing returns, with accuracy slightly declining. This suggests that while higher dimensions can capture more complex features, excessively high dimensions add unnecessary complexity without substantial performance gains.
Effect of Number of Stages. Increasing the number of stages in the model generally improves accuracy. The 4-stage configuration achieves the highest accuracy of 85.3% on the ModelNet40 dataset and 85.9% on the ScanObjectNN dataset, suggesting that deeper models are better equipped to capture complex spatial relationships and improve classification performance. Refer to Fig. 5 (c) for a detailed comparison of performance across different stage configurations.
Impact of Sigma ($\sigma$). The $\sigma$ parameter of the Gaussian kernel ($\sigma$) determines the degree of locality in feature aggregation. For the ModelNet40 dataset, $\sigma=\textbf{0.35}$ and $\sigma=\textbf{0.4}$ yield the highest accuracy, effectively preserving both local and broader spatial contexts. On the other hand, for the ScanObjectNN dataset, we find that $\sigma=\textbf{0.3}$ is the most effective, achieving the best accuracy by maintaining a balance between capturing fine-grained details and minimizing noise. Smaller values ($\sigma<0.3$) fail to aggregate sufficient local features, while larger values ($\sigma>0.4$) introduce excessive smoothing, negatively impacting performance. These trends are illustrated in Fig. 5 (d).